An isosceles right triangle is removed from each corner of a square piece of paper so that a rectangle remains. What is the lenght of a diagonal of the rectangle if the sum of the areas of the cut-off pieces is 200?
Answer by ankor@dixie-net.com(22740)

An isosceles right triangle is removed from each corner of a square piece of paper so that a rectangle remains. What is the length of a diagonal of the rectangle if the sum of the areas of the cut-off pieces is 200?
:
Each triangle has an area of 200/4 = 50: let x = length of the two equal sides of the triangle.
:
Find x:
.5(x^2) = 50
x^2 = 100
x = 10
:
The length of the side of the original square would be 2x or 20
:
The diagonal of the formed square would be the same, 20